OpenStudy (anonymous):

Need Help! Will give medal!!! A piece of wire 4 m long is cut into two pieces. One piece is bent into the shape of a circle of radius r and the other is bent into a square of side s. How should the wire be cut so that the total area enclosed is: a) a maximum? r= and s= . b) a minimum? r= and s= .

OpenStudy (pawanyadav):

For maximum, dA/dx=0 let parameter of square is x So area=x^2/16 Then 2πr=(4-x) r=4-x/2π Area=πr^2=(4-x)^2/4π Add area of square and circle and differentiate it to 0 ,then find x

OpenStudy (anonymous):

still a little lost

OpenStudy (anonymous):

I have no idea what you are talking about

OpenStudy (irishboy123):

fine :p do you want to walk through it? ie a solution?

OpenStudy (pawanyadav):

In which part

OpenStudy (anonymous):

both

OpenStudy (pawanyadav):

I think you are wanting a completecomplete solution

OpenStudy (irishboy123):

@kdancer03 @Pawanyadav split the wire into lengths x and y then used the x length to make a square and the y length to make a circle for the square, if it has perimeter x, it has area \(\dfrac{x^2}{16}\) ie \(\dfrac{x}{4} \times \dfrac{x}{4}\) good so far?

OpenStudy (anonymous):

that I get

OpenStudy (irishboy123):

the remaining bit of length y forms a circle of circumference y. and from \(C = 2 \pi r\) we know that that circle's radius can be gotten from \(y = 2 \pi r\) so \(r = \dfrac{y}{2 \pi}\) its area is \(A = \pi r^2 = \pi (\dfrac{y}{2 \pi})^2 = \dfrac{y^2}{4\pi}\) ok?

OpenStudy (anonymous):

yes

OpenStudy (irishboy123):

and because x + y = 4, we have x = 4-y and y = 4-x so we can convert the square area to a term in y, or the circle are to a term in x and then combine them make sense? or should we just do it?

OpenStudy (irishboy123):

recap \(A_{s}(x) = \dfrac{x^2}{16}\) \(A_{c}(y) = \dfrac{y^2}{4\pi}\) and \(x + y = 4\)

OpenStudy (anonymous):

I gotcha still

OpenStudy (irishboy123):

so now you decide do you go for A(y) or A(x), reducing this to single variable calculus. and then you set \(A' = 0\) to find a stationary point, ie max/min.....